OpenStudy (anonymous):

Solve using substitution: 3y − x − 1 = 0 - 4x = -6y A) No solution B) Infinitely many solutions C. (7, 8/3) D. (1, 2/3)

OpenStudy (anonymous):

How do you think it should be solved? It's just like the last two you posted in the last 10 min. Post how you think it should be solved and we can help you understand it.

OpenStudy (anonymous):

ans is D, i.e (x,y)=(1,2/3 )

OpenStudy (anonymous):

Find x from first equation : x = 3y - 1 Put it in second equation : -4(3y - 1) = -6y So, y = 2/3 And x = 1.. Answer is D...

OpenStudy (anonymous):

-4x=6y can be rewritten as x=6y-4 3y-(6y-4)=0 y = 2/3 plug back in, x =1